Solve:
2x + 8(x-3) = 4 - (3x +2)

Answer by checkley77(12844)

2x + 8(x-3) = 4 - (3x +2)
2x+8x-24=4-3x-2
10x+3x=24+4-2
13x=26
x=26/13
x=2 answer.
proof
2*2+8(2-3)=4-(3*2+2)
4+8*-1=4-(6+2)
4-8=4-8
-4=-4